The length of a rectangle is five times its width. If the perimeter of the rectangle is 108 m , find its area.
Answer by EMStelley(208)

Let's call the length L and the width W. Then since the length is five times the width, we have that L = 5W. Now, remembering that the perimeter of a rectangle is 2L+2W, we have
2L+2W=108
Now using the fact that L=5W we have:
2(5W)+2W=108
10W+2W=108
12W=108
W=9
So the width is 9m, which makes the length 9*5=45m.